Answer the following questions regarding this distribution: (1, 1, 2, 2, 3, 4, 4, 4, 5, 6, 7, 8} a. Is it symmetric or asymmetric? 2inlo Symmetric Asymmetric
Cognitive psychologist would like to evaluate the claim that the omega-3 fatty acids in fish can help improve memory in normal adult humans. One group of participants is given a pills containing a large dose of fish extract containing the Omega-3 (500 mg per day), and a second group is given a placebo (sugar pill). The researcher asks each participant to read the front page of the local newspaper thoroughly every day and to take one pill every day. The researcher gives each participant a memory test at the end of two weeks and records how many news items each participant remembers from the past three weeks of news. Answer the following questions about this scenario: a) What names would you give the independent and dependent variables? b) Is the dependent variable discrete or continuous? c)What scale of measurement (nominal, ordinal, interval or ratio; and continuous or discrete) is used to measure the independent variable? d) What research method is being used (experimental or observational)? Explain why you conclude that the research method is one or the other.
Answer #1

Problem 1:   

Min. 1st Qu. Median Mean 3rd Qu. Max.
1.000 2.000 4.000 3.917 5.250 8.000

Since 3rd Quartile-Median=5.250-4.000=1.250<Median-1st quartile=4-2=2

So the distribution is negatively skewed.

Hence Asymmetric.

Problem 2:

(a) Dependent variable: no. of news items remembered by a participant from the past three week

Independent variable: whether the participant takes Omega-3 or sugar pill.

(b) Discrete variable

(c) Nominal.

(d) Experimental. If the researcher has no control over the variables in an observational study whereas an experiment is a method of applying treatments to a group and recording the effects. Here we divide the participants into two groups and one group is received Omega-3 (i.e. treatment) and other group is received placebo (i.e. sugar pill) and then we see the effect on memory of a participant at the end of two weeks by taking the records how many news items each participant remembers from the past three weeks of news. Hence it is experimental.